OpenStudy (anonymous):

Please help me with this!!!! Find the slope‐intercept equation of the line with the following properties: Parallel to the line 5x + 2y = 1; containing the point (-5,4)

OpenStudy (anonymous):

2y=1-5x y=-5/2x+1/2 y=mx+b @Juliana90

OpenStudy (anonymous):

Is that the answer? Im really not very good at math, at all!

OpenStudy (anonymous):

Yes this is answer...

OpenStudy (anonymous):

Ok can you help me with one more?

OpenStudy (anonymous):

yes sure...

OpenStudy (anonymous):

Write the equation of the circle in standard form. Find the center, radius, intercepts, and graph the circle x^2+y^2-12x-4y+36=0

OpenStudy (anonymous):

shahzadjalbani, it seams, you have to deal with entire homework ...:)

OpenStudy (anonymous):

@Juliana90